Enough of a monoprotic acid is dissolved in water to produce a 1.15 M solution. The pH of the resulting solution is 2.99. Calculate the Ka for the acid.
Answer #1


use:
pH = -log [H+]
2.99 = -log [H+]
[H+] = 1.023*10^-3 M
HA dissociates as:

HA          ----->     H+   + A-
1.15                 0         0
1.15-x               x         x


Ka = [H+][A-]/[HA]
Ka = x*x/(c-x)
Ka = 1.023*10^-3*1.023*10^-3/(1.15-1.023*10^-3)
Ka = 9.114*10^-7
Answer: 9.11*10^-7